10 points are given on a plane. how many line segments can be drawn using them?
Answer by greenestamps(13200)



Any two points in a plane determine a line segment.

The number of different ways to choose 2 of the 10 given points is "10 choose 2", 10C2 = (10*9)/2 = 45.

So, as the problem is stated, the answer is 45.

Note the answer is different if you meant the number of LINES that could be drawn. In that case the answer would be 45 only if it is given that no 3 of the 10 points are collinear.

n collinear points determine only 1 line; but they determine nC2 = (n(n-1))/2 line segments.